### v1

# Necessary Conditions for Earthly Life Floating in the Venusian Atmosphere

Abstract
Millimeter-waveband spectra of Venus from both the James Clerk Maxwell Telescope (JCMT) and the Atacama Large Millimeter/submillimeter Array (ALMA) provide conclusive evidence (signal-to-noise ratio of about $15\sigma$) of a phosphine absorption-line profile against the thermal background from deeper, hotter layers of the atmosphere. Phosphine is an important biomarker; e.g., the trace of phosphine in the Earth’s atmosphere is uniquivocally associated with anthropogenic activity and microbial life (which produces this highly reducing gas even in an overall oxidizing environment). Motivated by the JCMT and ALMA tantalizing observations we reexamine whether Venus could accommodate Earthly life. More concretly, we hypothesize that the microorganisms populating the venusian atmosphere are not free floating but confined to the liquid environment inside cloud aerosols or droplets. Armed with this hypothesis, we generalize a study of airborne germ transmission to constrain the maximum size of droplets that could be floating in the venusian atmosphere and estimate whether their Stokes fallout times to reach moderately high temperatures are pronouncedly larger than the microbe’s replication time. We also comment on the effect of cosmic ray showers on the evolution of aerial microbial life.
I Introduction
It has long been known that the surface of Venus is too harsh an environment for life Sagan:1961 ; Sagan:1967 . Observations from Mariner 2 and Venus 4 spacecraft seem to indicate that the surface of Venus is a hot dielectric carrying temperatures of roughly $427^{\circ}{\rm C}$. The critical point of water is $374^{\circ}{\rm C}$ and 218 atmospheres, and hence liquid water at the average Venus surface temperature is out of the question. Furthermore, at temperatures well below $427^{\circ}{\rm C}$ enzymes are speedily inactivated, proteins denatured, and most biological organic molecules pyrolized. All in all, it seems quite safe to conjecture that the mean surface temperature of Venus excludes terrestrial forms of life.
Contrariwise, it has long been speculated that the clouds of Venus offer a favorable habitat for life, but regulated to be domiciled at an esentially fixed altitude Morowitz:1967 . The archetype living thing being the spherical hydrogen gasbag isopyenic organism (sHgio). The maximum sHgio size can be obtained by demanding that its mass must be equal to the displaced mass of the atmosphere, viz.,
$\rho_{\rm H}\ R_{2}^{3}+\rho_{\rm skin}\ (R_{1}^{3}-R_{2}^{3})=\rho_{\rm atm}\ R_{1}^{3}\,,$ (1)
where $\rho_{\rm atm}\sim 7\times 10^{-4}~{}{\rm g/cm}^{3}$ and $\rho_{\rm H}\sim 5\times 10^{-5}~{}{\rm g/cm}^{3}$ are respectively the atmospheric and hydrogen densities at 0.5 atm pressure level, and $\rho_{\rm skin}=1.1~{}{\rm g/cm}^{3}$ is the density of the skin modelled as a membrane of outer radius $R_{1}$ and inner radius $R_{2}$. For a minimum skin thickness of roughly $1~{}\mu{\rm m}$, i.e.,
$\frac{R_{1}-R_{2}}{R_{1}}\simeq 2\times 10^{-4}\,,$ (2)
Eq. (1) implies that sHgios would have a minimum diamter of $2R_{1}\sim 4~{}{\rm cm}$.
[…]
The JCMT and ALMA intriguing observations have reinvigorated investigations looking into the possibility of life in the atmosphere of Venus Bains ; Seager ; Clements . In particular, it was proposed in Seager that microbial life could reside inside liquid droplets/aerosols, which could protect the microbes from a fatal net loss of liquid to the atmosphere, an unavoidable problem for any free-floating living thing. However, the aerosol habitat could only have a limited lifetime because it would inexorably grow into droplets of a large enough size that are forced by gravity to settle downward to hotter, uninhabitable layers of the atmosphere. In this communication we generalize a study of airborne coronavirus transmission Anchordoqui to constrain the maximum size of the droplets that could be floating in the venusian atmosphere and estimate whether their Stokes fallout times to reach moderately high temperatures are pronouncedly smaller than the microbe’s replication time.
The layout of the paper is as follows. In Sec. II we provide a concise summary of the aerial life cycle put forward in Seager . In Sec. III we compare the Stokes fallout time of droplets with typical microbe’s replication times on Earth (under optimum conditions) to determine the maximum allowed radius for the life cycle to persist indefinitely in the atmosphere of Venus. In Sec. IV we comment on the effect of cosmic ray showers on the evolution of aerial microbial life. The paper wraps up in Sec. V with conclusions.
Figure 1: Hypothetical life cycle of the venusian microorganisms. Top panel: Cloud cover on Venus is permanent and continuous, with the middle and lower cloud layers at temperatures that are suitable for life. Bottom panel: Cycle for venusian aerial microbial life; see text for details. Taken from Ref. Seager .
II Life Cycle for Venusian Aerial Microbes
Venus’ clouds encircle the entire planet, resulting in a high planetary albedo $\sim 0.8$ Marov . The base of this thick cloud cover is situated at roughly 47 km above the surface (the temperature at this position is around $100^{\circ}$) and extends up to over 70 km in altitude. In equatorial and mid-latitudes the cloud top is located at 74 km, but decreases towards the poles to roughly 65 km Ignatiev . The size distribution of aerosol particles drifting inside the clouds enable a subdivission into three layers: upper (56.5 to 70 km altitude), middle (50.5 to 56.5 km), and lower (47.5 to 50.5 km); the smallest type-1 droplets (with a radius of $0.2~{}\mu{\rm m}$) and type-2 droplets (with a radius of 1 to 2 $\mu{\rm m}$) are present in all three cloud layers, whereas the largest type-3 droplets (with a radius of $4~{}\mu{\rm m}$) are only present in the middle and lower cloud layers Knollenberg .
It has long been suspected that the cloud decks of Venus offer an aqueous habitat where microorganisms can grow and flowrish Grinspoon . Carbon dioxide, sulfuric acid compounds, and ultraviolet (UV) light could give microbes food and energy. It is generally accepted that existence of life is most likely at an altitude of 50 km, where the temperature is between 60 and 90 degrees Celsius (140 and 194 degrees Fahrenheit), and the pressure is about 1 atm. An optimist might even imagine that the microbial life actually arose in a good-natured surface habitat, perhaps in a primitive ocean, before the planet suffered a runaway greenhouse, and the microbes lofted into the clouds Schulze-Makuch .
Even though the Earth’s atmosphere does not provide long-lived aerosols with temperatures or acidities comparable to Venus, the presence of life at high altitudes has been well-documented: bacteria, pollen, and algae have been observed in the Earth’s atmosphere as high as 15 km Womack . Furthermore, evidence has been found for the growth of bacteria in droplets sampled from a super-cooled cloud near a meteorological station on a mountain top in the Alps Sattler . It is commonly understood that these microorganisms would have likely reached such heights through evaporation, storms, eruptions, or meteor impacts. Of course, all of these processes could also have occurred in the early history of Venus. In contrast to the Earth, venusian clouds are not transient entities but represent a global, uninterrupted phenomenon, with the potential for aerosol particles to be sustained for a long period of time, rather than just a few days as in the terrestrial atmosphere. Thereby, the venusian clouds could provide a stable niche if microbes remain lofted in this aerial.
A cycle for venusian aerial microbial life was developed in Seager . The general idea, which is shown in Fig. 1, follows five steps that can be summarized as follows:
  1. 1.
    The cycle begins with dormant desiccated spores (black blobs in Fig. 1) which partially populate the lower haze layer of the atmosphere. [^1] In the spirit of Seager , throughout we use the term “spore” to indicate a cell in a dormant state of long-term metabolic inactivity, which is further resistant to (and protected from) environmental stresses.
  2. 2.
    Updraft of spores transports them up to the habitable layer. For example, the spores could travel up to the clouds via the effect of gravity waves. Despite the fact that gravity waves can only lead to the net transport of energy and momentum and not matter, they can compress atmosphere layers as they travel, producing vertical winds (which have been measured directly by the Venera landing probes 9 and 10 at the atmospheric lower haze layers Kerzhanovich ).
  3. 3.
    Shortly after reaching the (middle and lower cloud) habitable layer, the spores act as cloud condensation nuclei, and once surrounded by liquid (with necessary chemicals dissolved) germinate and become metabolically active.
  4. 4.
    Metabolically active microbes (dashed blobs in Fig. 1) grow and divide within liquid droplets (shown as solid circles in the figure). The liquid droplets grow by coagulation.
  5. 5.
    The droplets reach a size large enough to gravitationally settle down out of the atmosphere; higher temperatures and droplet evaporation trigger cell division and sporulation. The spores are small enough to withstand further downward sedimentation, remaining suspended in the lower haze layer (a depot of hibernating microbial life) to restart the cycle.
One of the key assumptions of the aerial life cycle put forward in Seager is the timescale on which droplets would persist in the habitable layer to empower replication. It is this that we now turn to study.
III Replication rates and Fallout times
Self-replication is a faculty inherent to every species of living thing, and standard intuition imposes that such a physical process must invariably be fueled by the production of entropy Chudnovsky:1985 . It is of interest then to estimate a lower bound for the amount of heat that is produced during a process of self-replication in a system coupled to a thermal bath. The minimum value for the physically allowed rate of heat production is determined by the growth rate, internal entropy, and durability of the replicator. Impressively, bacteria replicate close (within a factor of two or three) to the physical efficiency England:2012 .
Bacteria replicate by binary fission, a process by which one bacterium splits into two. Therefore, bacteria increase their numbers by geometric progression whereby their population doubles every generation time. Generation time is the time it takes for a population of bacteria to double in number. For many common bacteria, the generation time is quite short, 20 to 60 minutes under optimum conditions. The typical example is Escherichia coli (or E. coli), which can divide every 20 minutes under aerobic, nutrient-rich conditions (but of course the generation time for bacteria in the wild are substantially greater than those in the laboratory) Gibson . Actually, Vibrio natriegens (previously known as Pseudomonas natriegens and Beneckea natriegens) is a free-living marine bacterium with the fastest generation time known, between 7 to 10 minutes Eagon ; Maida ; Lee . In our calculations we adopt as benchmark the E. coli generation time under optimum conditions. The duplication times of other types of bacterium are listed in Table 1.
Table 1: Bacteria generation times.
Name Reproduction Time Ref.
Bacillus subtilis 20 minutes Errington
Pseudomonas aeruginosa 16 to 24 hours LaBauve
Vibrio cholerae 20 minutes Das
Bacillus thuringiensis 20 minutes Choi
Shigella flexneri 40 minutes Jennison
Streptococcus pyogenes 12 to 16 hours Gera
Salmonella typhimurium 20 minutes Lowrie
Clostridium perfringens          10 to 12 minutes Li
Pseudomonas fluorescens 1.5 hours Caldwell
The mean division time for bacteria population is then 20 minutes. If the observation begins with one bacterium, we can estiamte how many bacteria will be present after six hours. The E. coli divides every 20 minutes, and so this bacterium divide (60/20 =3) three times every hour. If the bacteria grow for twelve hours, each bacterium will divide 3 times per hour $\times$ 12 hours = 36 times. Every time the bacteria reproduce, the number doubles. Then, the number of bacteria at the end of the growth period is found to be
[…]
where $N_{\rm initial}$ is number of bacteria at the beginning of the growth period and $n$ is the number of divisions. For $N_{\rm initial}=1$ and $n=36$ we have $N_{\rm final}\sim 7\times 10^{10}$ bacteria. We conclude that to mantain the colony of microbes alive we require Stokes fallout times longer than half an Earth day.
Assuming that the droplets are spherical the mass can be simply estimated as
$m_{\rm droplet}=\frac{4}{3}\pi r_{\rm droplet}^{3}\ \rho\,,$ (4)
where $\rho$ is the droplet’s density; in our calculations we consider the E. coli dry mass density $\rho_{\rm Ec}\sim 300~{}{\rm kg/m}^{3}$ Pang and the density of water $\rho_{\rm H_{2}O}\sim 997~{}{\rm kg/m}^{3}$.
Under the action of gravity, droplets of mass $m_{\rm droplet}$ and size $r_{\rm droplet}$ would acquire a downward terminal speed that follows from Stokes law and is given by
$v_{\rm down}=\mu\ m_{\rm droplet}\ g_{{}_{\venus}}\,,$ (5)
where $g_{{}_{\venus}}$ is the acceleration due to gravity in Venus,
$\mu=\frac{1}{6\pi\ \eta_{{}_{\venus}}\ r_{\rm droplet}}$ (6)
is the organism mobility in the fluid, and where $\eta_{{}_{\venus}}$ is the dynamic viscosity of the venusian atmosphere Einarsson . For the dynamic viscocity, we adopt the estimate of Ref. Beck , $\eta_{{}_{\venus}}\sim 2\times 10^{-5}~{}{\rm kg/m/s}$.
Figure 2: Terminal speed as a function of the microbe’s radius.
The acceleration of gravity at the surface of Venus can be derived from measurements of Mariner V Anderson:1967
$G_{N}M_{{}_{\venus}}=324,856.6\pm 0.5~{}{\rm km^{3}/s^{2}}\,,$ (7)
and is found to be $g_{{}_{\venus}}\simeq 8.9~{}{\rm m/s^{2}}$, where $G_{N}$ is Newton’s gravitational constant and where we have taken the radius of Venus to be $R_{{}_{\venus}}\simeq 6,056~{}{\rm km}$ to accommodate existing observations Ash:1967 ; Ash:1968 ; Anderson:1968 .
Putting all this together, we can now combine Eqs. (5) and (6) to estimate the downward terminal speed of the droplets. In Fig. 2 we show the terminal velocity as a function of the droplet’s radius. As one can check by inspection, the terminal speed of aerosols (viz. droplets with $r_{\rm droplet}\lesssim 1~{}\mu{\rm m}$) is negligible, and so we can conclude that gravity would play no role in the motion of the microbes through the atmosphere. As the droplet size approaches $100~{}\mu{\rm m}$ the droplets would start sinking to the lower haze layers.
IV Cosmic ray effects on microbial life
We have long been suspecting that cosmic radiation played a pivotal role in the evolution of life on Earth Joly ; for a more recent prespective see Dartnell . In this section, we briefly comment on whether energetic particle radiation could provide a threat for any type of microbe populating the clouds of Venus.
Pioneer Venus measurements are consistent with the absence of an intrinsic magnetic $\vec{B}$ field that could provide a shield against low and high energy charged particles reaching the atmposphere Russell . If taken at face value, the lack of $\vec{B}$ field seems to indicate that the solar wind would interact directly with the upper atmosphere. However, the upper atmosphere is ionized by UV radiation, yielding the so-called “ionosphere.” Currents arising from the interaction between the solar wind and the electrically conductive venusian ionosphere induce a magnetic field, and so the incoming solar wind particles are slowed down and diverted around the planet Zhang . However, the induced magnetic field is weak Russell:2006 and can only deflect charged energetic particles with energies up to several hundreds of keV. Thus, most of the cosmic radiation have unrestricted access to the Venusian atmosphere.
Cosmic rays with energies above about 1 GeV induce particle cascades via inelastic scattering with atmospheric nuclei, producing fluxes of secondary, tertiary, and subsequent generations of particles. All these particles together create a cascade (or shower). As the cascade develops longitudinally the particles become less and less energetic since the energy of the incoming cosmic ray is redistributed among more and more participants. The transverse momenta acquired by the secondaries cause the particles to spread laterally as they propagate through the atmospheric target Anchordoqui:2018qom .
As can be seen in Fig. 1, the top region of the temperate zone is at 62 km altitude, and thus it receives less than $200~{}{\rm g/cm^{2}}$ shielding depth against cosmic ray showers. This is far less than the biosphere on the surface of the Earth, which is beneath $1033~{}{\rm g/cm^{2}}$. The atmospheric grammage at the middle cloud layer is, however, $1000~{}{\rm g/cm^{2}}$, suggesting the cosmic ray induced effects would be comparable to those on the Earth’s surface. Indeed, numerical simulations show that cosmic radiation would not have had any hazardous effect on putative microorganisms within the potentially temperate zone (51 to $62~{}{\rm km}$Dartnell:2015 ; Herbst .
V Conclusions
We have re-examined whether one can envisage life inside the clouds of Venus, operating entirely on known terrestrial principles. We have shown that for aerosols, the Stokes fallout times to reach the lower haze atmospheric layers is pronouncedly larger than the typical bacterium replication time on Earth. Bearing this in mind, we can conclude that if updraughts exist, a stable population of microorganisms that in the early history of Venus emigrated from the surface to the atmospheric clouds and now remain confined to aerosols may be possible. Such an aerial microbial life may be awaiting for the Venus Life Finder (VLF) Missions to arrive Seager:2021m .
Acknowledgments
We thank William Bains and Janusz Petkowski for permission to reproduce Fig. 1. The research of LAA is supported by the U.S. NSF grant PHY-2112527.